Find three consecutive even integers such that three times the middle integers is four more than the sum of the first and third?
The sum of two number is eighteen. The total of the three times the smaller is equal to the larger. Find the two numbers.

Answer by stanbon(75887)

Find three consecutive even integers such that three times the middle integer is four more than the sum of the first and third?
----
1st: 2x-2
2nd: 2x
3rd: 2x+2
------
Equation:
3(2x) = 4x+4
6x = 4x+4
2x = 4
x = 2
----
1st: 2x-2 = 2
2nd: 4
3rd: 6
================
The sum of two number is eighteen. The total of the three times the smaller is equal to the larger. Find the two numbers.
---
Equations:
x + y = 18
3x = y
---
Substitute and solve for "x":
x+3x = 18
4x = 18
x = 9/2
-----
Solve for "y":
y = 18-x
y = (36/2)-(9/2)
y = 27/2
------
Cheers,
Stan H.
==========